Given the function below, algebraically solve the following system of equation. Please show all work.


y= x^2+2x+3

2y-2x=10


1
Expert's answer
2021-03-02T05:11:41-0500

2*y-2*x=10 is the same that y - x = 5 "\implies" y = x + 5. Substitute this to first expression:

y = x2+2x+3 "\implies" x + 5 = x2+2x+3 "\implies" x2+ x - 2 = 0 "\implies" x1 = 1 , x2 = -2. Hence y1 = 6, y2 = 3.